# Compact manifolds of dimension $n\geq 12$ with positive isotropic curvature

### v21

@@ 2 Orbifold ancient $\kappa$-solutions and standard solutions
Proof of Claim 1. Given any $2\varepsilon_{4}$-neck $N$ in $(\mathcal{O},g)$, note that the diameter of $\pi(\mathbb{S}^{n-1}\times\{0\})$ is much smaller than that of the neck $N$ since $\varepsilon_{4}$ is very small. It follows that at least four ninths of the neck $N$ is contained in a standard cylindrical neck in $\mathcal{O}$. By Proposition 5.4 any cross section of the neck $N$ is diffeomorphic to $\mathbb{S}^{n-1}$, so the neck $N$ is diffeomorphic to $\mathbb{S}^{n-1}\times(0,1)$. It is well-known that for $n\geq 5$ (in fact for all $n\geq 3$), $\mathbb{R}P^{n-1}$ cannot be embedded in $\mathbb{R}^{n}$, in particular, $\mathbb{R}P^{n-1}$ cannot be embedded in $\mathbb{S}^{n-1}\times(0,1)$. So the intersection of the central cross section $\Sigma$ of the neck $N$ with $\pi(\mathbb{S}^{n-1}\times\{0\})$ is empty. It follows that the middle half of the neck $N$ is contained in a standard cylindrical neck, say $N^{\prime}$, in $\mathcal{O}$. By Proposition 5.4 again we see that the conclusion of Claim 1 holds. (In fact, by Proposition 5.4 we see that the compact domain $\Omega$ bounded by $\Sigma$ is diffeomorphic to $\mathbb{R}P^{n}\setminus B$, as any cross section of a standard cylindrical neck in $(\mathcal{O},g)$ bounds such a domain.) By the way, once we know that the neck $N$ is diffeomorphic to $\mathbb{S}^{n-1}\times(0,1)$ and the middle half of the neck $N$ is contained in a standard cylindrical neck, we can also proceed as follows. Recall the following result: Let $n\geq 5$. If $S$ is a smoothly embedded $(n-1)$-sphere in $D^{n}\setminus\partial D^{n}$, then $S$ bounds a smoothly embedded $n$-disk $D^{\prime}$; moreover, the complement of $\text{Int}\hskip 1.42262ptD^{\prime}$ in $D^{n}$ is diffeomorphic to $\mathbb{S}^{n-1}\times[0,1]$. (The former fact follows from the solution of Schoenflies conjecture in dimensions $n\geq 5$, and the latter fact follows from the former fact and the Cerf-Palais disk theorem.) Then arguing as in the proof of Corollary A.2.2 and Lemma 3.2.2 in [6], we see that if $\varepsilon_{4}$ is sufficiently small, $\Sigma$ and certain cross section of $N^{\prime}$ cobounds a compact domain which is diffeomorphic to $\mathbb{S}^{n-1}\times[0,1]$, and Claim 1 follows. $\Box$
@@
Proof of Claim 2. Given any $2\varepsilon_{4}$-neck $N$ in $(\mathcal{O},g)$, as noted above, the two singular points of $\mathcal{O}$ are contained in $\pi(\mathbb{S}^{n-1}\times\{0\})$, whose diameter is much smaller than that of the neck $N$ if $\varepsilon_{4}$ is sufficiently small. Of course none of the two singular points is contained in $N$. So the intersection of the central cross section $\Sigma$ of the neck $N$ with $\pi(\mathbb{S}^{n-1}\times\{0\})$ is empty. It follows that the middle half of the neck $N$ is contained in a standard cylindrical neck, say $N^{\prime}$, in $\mathcal{O}$. Now Claim 2 follows from Proposition 5.4. $\Box$
@@
where $Q$ is a positive constant and $g_{\text{cyl}}$ is the standard metric on the cylinder $\mathbb{S}^{n-1}\times(-(2\varepsilon_{4})^{-1},(2\varepsilon_{4})^{-1})$ with scalar curvature 1 and length ${\varepsilon_{4}}^{-1}$ for the interval $(-(2\varepsilon_{4})^{-1},(2\varepsilon_{4})^{-1})$. Suppose that the image of $\mathbb{S}^{n-1}\times(-\frac{1}{2\varepsilon_{4}},0]$ under the map $\psi$ is contained in $\Omega$. Following [20], [48], and [55], let $\varphi:(0,B]\rightarrow\mathbb{R}$ be a smooth function as chosen in the appendix, where $B>\frac{1}{10}$; in particular, $\varphi(z)=ce^{-1/z}$ for $z\in(0,\frac{1}{10}]$, where $c<\frac{1}{\sqrt{(n-1)(n-2)}}$ is a positive constant. (We require that $\frac{1}{\varepsilon_{4}}\gg B$.) We also choose a smooth cutoff function $\chi:(\frac{1}{20},B]\rightarrow[0,1]$ with $\chi(z)=1$ for $\frac{1}{20}<z\leq\frac{1}{18}$ and $\chi(z)=0$ for $\frac{1}{12}\leq z\leq B$. For convenience, below for $x\in\psi(\mathbb{S}^{n-1}\times(0,B])$ we will also denote $z(x)=p_{2}\circ\psi^{-1}(x)$ by $z$, where $p_{2}:\mathbb{S}^{n-1}\times(-(2\varepsilon_{4})^{-1},(2\varepsilon_{4})^{-1})\rightarrow(-(2\varepsilon_{4})^{-1},(2\varepsilon_{4})^{-1})$ is the projection onto the second factor. Let
$g^{\prime}(x)=\begin{cases}g(x),&{x\in\Omega;}\\ e^{-2\varphi(z)}g(x),&{z\in(0,\frac{1}{20}];}\\ e^{-2\varphi(z)}(\chi(z)g(x)+(1-\chi(z))Q^{-1}(\psi_{*}g_{\text{cyl}})(x)),&{z\in(\frac{1}{20},B],}\end{cases}$
@@
which gives a new metric on the orbifold $\Omega\cup\psi(\mathbb{S}^{n-1}\times(0,B])$. If $\varepsilon_{4}$ is sufficiently small, the new metric $g^{\prime}$ will remain to be weakly PIC2 and strictly PIC. Moreover, as in Section 72 of [55] and the appendix, $\varphi$ is chosen such that the metric $g^{\prime}$ restricted to the portion $\psi(\mathbb{S}^{n-1}\times[B^{\prime},B])$ has constant positive sectional curvature, where $\frac{1}{10}<B^{\prime}<B$ is as chosen in the appendix. We extend the metric $g^{\prime}$ by smoothly attaching a metric $n$-ball of constant positive sectional curvature to $(\Omega\cup\psi(\mathbb{S}^{n-1}\times(0,B]),g^{\prime})$ along the boundary such that the extended metric is weakly PIC2 and strictly PIC; moreover, the Ricci curvature (even sectional curvature) of the extended metric is positive at (at least) one point. We denote the resulting closed Riemannian orbifold by $(\mathcal{O}^{\prime},g^{\prime})$. For more details see also the proof of Proposition 8.2 in [20] (but note that in the definition of the function $\varphi$ there, one should add a small positive factor (depending on the dimension) before $e^{-\frac{1}{z}}$), Section 72 of [55], Chapter 7 of [6], and Chapter 13 of [67]. By the way, it seems that the formula for the function $f$ near the point $z=4$ displayed on p. 424 of [27] and p. 228 of [31] (and p. 1221 of [50] and p. 343 of [83]) is not quite true; cf. Lemma 2.10 in [36] and Chapter 1 in [73].
@@
We can write $\mathcal{O}^{\prime}=D_{1}\cup\Omega$, where $D_{1}$ is diffeomorphic to $D^{n}$, and $D_{1}\cap\Omega=\partial D_{1}=\partial\Omega=\Sigma$. By inspecting the surgery procedure we see that $f$ extends to a diffeomorphism from $D^{n}$ to $D_{1}$, which factors through a diffeomorphism from $D^{n}$ to a rotationally symmetric surgery cap constructed in the appendix. Then we run the orbifold Ricci flow with initial data $(\mathcal{O}^{\prime},g^{\prime})$. By Proposition 2.4 this orbifold Ricci flow solution is strictly PIC2 for any $t>0$. By the differentiable sphere theorem for compact orbifolds which are strictly PIC2 we see that $(\mathcal{O}^{\prime},g^{\prime})$ is diffeomorphic to a spherical orbifold with two singular points both with local group $\mathbb{Z}_{2}$, which must be of the form $\mathbb{S}^{n}//(x,\pm x^{\prime})$ by using Lemma 2.2 above, Lemma 5.2 in [29] and Corollary 2.4 in Chapter VI of [15]. Note that one can extend the differentiable sphere theorem for compact manifolds which is strictly PIC2 in [23] to the orbifold case either by using the estimates in [23], Perelman’s noncollapsing theorem [71] (which can be easily extended to the orbifold case) and the orbifold Ricci flow compactness theorem [56][62], or by first running the orbifold Ricci flow which is strictly PIC2 up to certain time such that the sectional curvature is globally strictly quarter pinched (the existence of such time is guaranteed by the estimates in [23]), then using Proposition 5.2 in Böhm-Wilking [8], which will reduce the orbifold case to the manifold case. Now we have a diffeomorphism $H:\mathbb{S}^{n}//(x,\pm x^{\prime})=\bar{B}\cup(\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus B)\rightarrow\mathcal{O}^{\prime}=D_{1}\cup\Omega$. Note that both $f\circ f^{\prime}$ and $f\circ r_{1}\circ f^{\prime}$ extend to a diffeomorphism from $\bar{B}$ to $D_{1}$. We choose one of the extensions such that this extension and $H|_{\bar{B}}$ are either both orientation-preserving or both orientation-reversing if $n$ is even. Then the desired result in this subcase follows from Lemma 2.3. $\Box$
@@
Definition. (cf. Definition 6.16 in [20]) Let $\varepsilon_{0}=\varepsilon_{0}(n)$ be a small positive constant and $0<\varepsilon<\frac{1}{4}\varepsilon_{0}$. Given a point $x_{0}\in(\mathcal{O},g)$, an open subset $U$ of $\mathcal{O}$ is an $\varepsilon$-cap centered at $x_{0}$ if $U$ is a topological cap and $U\setminus V$ is an $\varepsilon$-neck, where $V$ is a compact domain in $U$ diffeomorphic to some $D^{n}//\Gamma$ or $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus B$ or $\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus B$ with $\partial V$ a central cross section of an $\varepsilon$-neck (that is, the image of $\mathbb{S}^{n-1}/\Gamma\times\{0\}$ under the diffeomorphism associated to the $\varepsilon$-neck) and with $x_{0}\in\text{Int}\hskip 1.42262ptV$, (we will call $V$ a core of the cap $U$,) and in addition,
@@
The idea behind this definition is that an $\varepsilon$-cap $U$ should be, after suitable rescaling, $\eta$-close to the corresponding subset $W$ of an orbifold ancient $\kappa$-solution or a standard solution with an $\eta$-homothety $\psi:W\rightarrow U$, where $\eta$ is much smaller than $\varepsilon$ (compare p. 522 of [20] and Definition 69.1 in [55]); moreover, the definition should help one to identify some components which are discarded in the process of the surgery of the Ricci flow (see Proposition 3.2). Assume that the end of the $\varepsilon$-cap $U$ is diffeomorphic to $\mathbb{S}^{n-1}\times(0,1)$. Let $N_{1}$ be an $\varepsilon_{0}$-neck given by $\psi_{1}:\mathbb{S}^{n-1}\times(-\varepsilon_{0}^{-1},\varepsilon_{0}^{-1})\rightarrow N_{1}$ which is contained in the $\varepsilon$-neck end in $U$. (Note that by Proposition 5.4, given $0<\varepsilon<\varepsilon^{\prime}\leq 2\varepsilon_{1}$ and an $\varepsilon$-neck diffeomorphic to $\mathbb{S}^{n-1}\times(0,1)$, any $\varepsilon^{\prime}$-neck contained in this $\varepsilon$-neck is also diffeomorphic to $\mathbb{S}^{n-1}\times(0,1)$.) If $\varepsilon_{0}\leq\varepsilon_{4}$, by Lemmas 5.2 and 5.3, a suitable restriction of $\psi^{-1}\circ\psi_{1}$ will give a $2\varepsilon_{4}$-neck in $W$. Then we can use Proposition 2.5 or Proposition 2.6 to get the desired extra property for the $\varepsilon$-cap $U$.
@@
(cf. Theorem 6.18 in [20] and Theorem 3.9 in [29])   Given a small positive constant $\varepsilon$ and a constant $\theta>0$, there exist positive constants $C_{1}=C_{1}(n,\theta,\varepsilon)$ and $C_{2}=C_{2}(n,\theta,\varepsilon)$, such that given any noncompact orbifold ancient $\kappa$-solution $(\mathcal{O},g(t))$ of dimension $n\geq 5$ (with at most isolated singularities) which satisfies $Rm-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in PIC$ and is not locally isometric to an evolving shrinking round cylinder, for each space-time point $(x_{0},t_{0})$, there is an open subset $U$ with $B(x_{0},t_{0},C_{1}^{-1}R(x_{0},t_{0})^{-\frac{1}{2}})\subset U\subset B(x_{0},t_{0},C_{1}R(x_{0},t_{0})^{-\frac{1}{2}})$, which falls into one of the following categories:
(a) $U$ is an $\varepsilon$-neck centered at $(x_{0},t_{0})$,
(b) $U$ is an $\varepsilon$-cap centered at $(x_{0},t_{0})$ and diffeomorphic to $\mathbb{R}^{n}//\Gamma$ for some finite subgroup $\Gamma$ of $O(n)$ acting freely on $\mathbb{S}^{n-1}$;
@@
moreover, the scalar curvature in $U$ at time $t_{0}$ is (strictly) between $C_{2}^{-1}R(x_{0},t_{0})$ and $C_{2}R(x_{0},t_{0})$.

### v22

@@ 2 Orbifold ancient $\kappa$-solutions and standard solutions
Proof of Claim 1. Given any $2\varepsilon_{4}$-neck $N$ in $(\mathcal{O},g)$, note that the diameter of $\pi(\mathbb{S}^{n-1}\times\{0\})$ is much smaller than that of the neck $N$ since $\varepsilon_{4}$ is very small. It follows that at least four ninths of the neck $N$ is contained in a standard cylindrical neck in $\mathcal{O}$. By Proposition A.4 any cross section of the neck $N$ is diffeomorphic to $\mathbb{S}^{n-1}$, so the neck $N$ is diffeomorphic to $\mathbb{S}^{n-1}\times(0,1)$. It is well-known that for $n\geq 5$ (in fact for all $n\geq 3$), $\mathbb{R}P^{n-1}$ cannot be embedded in $\mathbb{R}^{n}$, in particular, $\mathbb{R}P^{n-1}$ cannot be embedded in $\mathbb{S}^{n-1}\times(0,1)$. So the intersection of the central cross section $\Sigma$ of the neck $N$ with $\pi(\mathbb{S}^{n-1}\times\{0\})$ is empty. It follows that the middle half of the neck $N$ is contained in a standard cylindrical neck, say $N^{\prime}$, in $\mathcal{O}$. By Proposition A.4 again we see that the conclusion of Claim 1 holds. (In fact, by Proposition A.4 we see that the compact domain $\Omega$ bounded by $\Sigma$ is diffeomorphic to $\mathbb{R}P^{n}\setminus B$, as any cross section of a standard cylindrical neck in $(\mathcal{O},g)$ bounds such a domain.) By the way, once we know that the neck $N$ is diffeomorphic to $\mathbb{S}^{n-1}\times(0,1)$ and the middle half of the neck $N$ is contained in a standard cylindrical neck, we can also proceed as follows. Recall the following result: Let $n\geq 5$. If $S$ is a smoothly embedded $(n-1)$-sphere in $D^{n}\setminus\partial D^{n}$, then $S$ bounds a smoothly embedded $n$-disk $D^{\prime}$; moreover, the complement of $\text{Int}\hskip 1.42262ptD^{\prime}$ in $D^{n}$ is diffeomorphic to $\mathbb{S}^{n-1}\times[0,1]$. (The former fact follows from the solution of Schoenflies conjecture in dimensions $n\geq 5$, and the latter fact follows from the former fact and the Cerf-Palais disk theorem.) Then arguing as in the proof of Corollary A.2.2 and Lemma 3.2.2 in [6], we see that if $\varepsilon_{4}$ is sufficiently small, $\Sigma$ and certain cross section of $N^{\prime}$ cobounds a compact domain which is diffeomorphic to $\mathbb{S}^{n-1}\times[0,1]$, and Claim 1 follows. $\Box$
@@
Proof of Claim 2. Given any $2\varepsilon_{4}$-neck $N$ in $(\mathcal{O},g)$, as noted above, the two singular points of $\mathcal{O}$ are contained in $\pi(\mathbb{S}^{n-1}\times\{0\})$, whose diameter is much smaller than that of the neck $N$ if $\varepsilon_{4}$ is sufficiently small. Of course none of the two singular points is contained in $N$. So the intersection of the central cross section $\Sigma$ of the neck $N$ with $\pi(\mathbb{S}^{n-1}\times\{0\})$ is empty. It follows that the middle half of the neck $N$ is contained in a standard cylindrical neck, say $N^{\prime}$, in $\mathcal{O}$. Now Claim 2 follows from Proposition A.4. $\Box$
@@
where $Q$ is a positive constant and $g_{\text{cyl}}$ is the standard metric on the cylinder $\mathbb{S}^{n-1}\times(-(2\varepsilon_{4})^{-1},(2\varepsilon_{4})^{-1})$ with scalar curvature 1 and length ${\varepsilon_{4}}^{-1}$ for the interval $(-(2\varepsilon_{4})^{-1},(2\varepsilon_{4})^{-1})$. Suppose that the image of $\mathbb{S}^{n-1}\times(-\frac{1}{2\varepsilon_{4}},0]$ under the map $\psi$ is contained in $\Omega$. Following [21], [51], and [58], let $\varphi:(0,b_{0}]\rightarrow\mathbb{R}$ be a smooth function as chosen in the appendix, where $b_{0}>\frac{1}{10}$; in particular, $\varphi(z)=ce^{-1/z}$ for $z\in(0,\frac{1}{10}]$, where $c<\frac{1}{\sqrt{(n-1)(n-2)}}$ is a small positive constant. (We require that $\frac{1}{\varepsilon_{4}}\gg b_{0}$.) We also choose a smooth cutoff function $\chi:(\frac{1}{20},b_{0}]\rightarrow[0,1]$ with $\chi(z)=1$ for $\frac{1}{20}<z\leq\frac{1}{18}$ and $\chi(z)=0$ for $\frac{1}{12}\leq z\leq b_{0}$. For convenience, below for $x\in\psi(\mathbb{S}^{n-1}\times(0,b_{0}])$ we will also denote $z(x)=p_{2}\circ\psi^{-1}(x)$ by $z$, where $p_{2}:\mathbb{S}^{n-1}\times(-(2\varepsilon_{4})^{-1},(2\varepsilon_{4})^{-1})\rightarrow(-(2\varepsilon_{4})^{-1},(2\varepsilon_{4})^{-1})$ is the projection onto the second factor. Let
$g^{\prime}(x)=\begin{cases}g(x),&{x\in\Omega;}\\ e^{-2\varphi(z)}g(x),&{z\in(0,\frac{1}{20}];}\\ e^{-2\varphi(z)}(\chi(z)g(x)+(1-\chi(z))Q^{-1}(\psi_{*}g_{\text{cyl}})(x)),&{z\in(\frac{1}{20},b_{0}],}\end{cases}$
@@
which gives a new metric on the orbifold $\Omega\cup\psi(\mathbb{S}^{n-1}\times(0,b_{0}])$. If $\varepsilon_{4}$ is sufficiently small, the new metric $g^{\prime}$ will remain to be weakly PIC2 and strictly PIC. Moreover, as in Section 72 of [58] and the appendix, $\varphi$ is chosen such that the metric $g^{\prime}$ restricted to the portion $\psi(\mathbb{S}^{n-1}\times[b^{\prime},b_{0}])$ has constant positive sectional curvature, where $\frac{1}{10}<b^{\prime}<b_{0}$ is as chosen in the appendix. We extend the metric $g^{\prime}$ by smoothly attaching a closed metric $n$-ball of constant positive sectional curvature to $(\Omega\cup\psi(\mathbb{S}^{n-1}\times(0,b_{0}]),g^{\prime})$ along the boundary such that the extended metric is weakly PIC2 and strictly PIC; moreover, the Ricci curvature (even sectional curvature) of the extended metric is positive at (at least) one point. We denote the resulting closed Riemannian orbifold by $(\mathcal{O}^{\prime},g^{\prime})$. For more details see also the proof of Proposition 8.2 in [21] (but note that in the definition of the function $\varphi$ there, one should add a small positive factor (depending on the dimension) in front of $e^{-\frac{1}{z}}$), Section 72 of [58], Chapter 7 of [6], and Chapter 13 of [70]. By the way, it seems that the formula for the function $f$ near the point $z=4$ displayed on p. 424 of [29] and p. 228 of [33] (and p. 1221 of [53] and p. 343 of [90]) is not quite true; cf. Lemma 2.10 in [39] and Chapter 1 in [77].
@@
We can write $\mathcal{O}^{\prime}=D_{1}\cup\Omega$, where $D_{1}$ is diffeomorphic to $D^{n}$, and $D_{1}\cap\Omega=\partial D_{1}=\partial\Omega=\Sigma$. By inspecting the surgery procedure we see that $f$ extends to a diffeomorphism from $D^{n}$ to $D_{1}$, which factors through an $O(n)$-equivariant diffeomorphism from $D^{n}$ to a rotationally symmetric surgery cap constructed in the appendix. Then we run the orbifold Ricci flow with initial data $(\mathcal{O}^{\prime},g^{\prime})$. By Proposition 2.4 this orbifold Ricci flow solution is strictly PIC2 for any $t>0$. By the differentiable sphere theorem for compact orbifolds which are strictly PIC2 we see that $(\mathcal{O}^{\prime},g^{\prime})$ is diffeomorphic to a spherical orbifold with two singular points both with local group $\mathbb{Z}_{2}$, which must be of the form $\mathbb{S}^{n}//(x,\pm x^{\prime})$ by using Lemma 2.2 above, Lemma 5.2 in [31] and Corollary 2.4 in Chapter VI of [16]. Note that one can extend the differentiable sphere theorem for compact manifolds which is strictly PIC2 in [24] to the orbifold case either by using the estimates in [24], Perelman’s noncollapsing theorem [75] (which can be easily extended to the orbifold case) and the orbifold Ricci flow compactness theorem [59][65], or by first running the (compact) orbifold Ricci flow which is strictly PIC2 up to certain time such that the sectional curvature is globally strictly quarter pinched (the existence of such time is guaranteed by the estimates in [24] and arguments in the proofs of Theorem 5.1 in [8] and Corollary 2.3.4 in [89]), then using Proposition 5.2 in Böhm-Wilking [8], which will reduce the orbifold case to the manifold case. Now we have a diffeomorphism $H:\mathbb{S}^{n}//(x,\pm x^{\prime})=\bar{B}\cup(\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus B)\rightarrow\mathcal{O}^{\prime}=D_{1}\cup\Omega$. Note that both $f\circ f^{\prime}$ and $f\circ r_{1}\circ f^{\prime}$ extend to a diffeomorphism from $\bar{B}$ to $D_{1}$. We choose one of the extensions such that this extension and $H|_{\bar{B}}$ are either both orientation-preserving or both orientation-reversing if $n$ is even. Then the desired result in this subcase follows from Lemma 2.3. $\Box$
Remark. By the way, one can show that if $G$ is a nontrivial, finite subgroup of $\text{Isom}(\mathbb{S}^{2k}\times\mathbb{R}^{l})$ acting freely on $\mathbb{S}^{2k}\times\mathbb{R}^{l}$, where $k$ and $l$ are positive integers, then $G\cong\mathbb{Z}_{2}$. We can argue as follows. Recall that $\text{Isom}(\mathbb{S}^{2k}\times\mathbb{R}^{l})\cong\text{Isom}(\mathbb{S}^{2k})\times\text{Isom}(\mathbb{R}^{l})$; see for example [83]. Suppose that $g\in G$ is nontrivial. As $g$ is of finite order, the $\text{Isom}(\mathbb{R}^{l})$-component of $g$ must have a fixed point in $\mathbb{R}^{l}$; let $q$ be such a fixed point. By restricting the action of $g$ to the slice $\mathbb{S}^{2k}\times\{q\}$ we see that the $\text{Isom}(\mathbb{S}^{2k})$-component of $g$ must be the antipodal map on $\mathbb{S}^{2k}$. Since the $\text{Isom}(\mathbb{R}^{l})$-component of $g^{2}$ must have a fixed point in $\mathbb{R}^{l}$, we see that $g^{2}=1$. Similarly, if $g_{1},g_{2}\in G$ are both nontrivial, then the $\text{Isom}(\mathbb{S}^{2k})$-component of $g_{i}$ must be the antipodal map on $\mathbb{S}^{2k}$, $i=1,2$, and the $\text{Isom}(\mathbb{S}^{2k})$-component of the product $g_{1}g_{2}$ must be trivial, but the $\text{Isom}(\mathbb{R}^{l})$-component of $g_{1}g_{2}$ has a fixed point in $\mathbb{R}^{l}$, so $g_{1}g_{2}=1$. As $g_{1}^{2}=1$ we see that $g_{1}=g_{2}$.
@@
Definition. (cf. Definition 6.16 in [21]) Let $\varepsilon_{0}=\varepsilon_{0}(n)$ be a small positive constant and $0<\varepsilon<\frac{1}{4}\varepsilon_{0}$. Given a point $x_{0}\in(\mathcal{O},g)$, an open subset $U$ of $\mathcal{O}$ is an $\varepsilon$-cap centered at $x_{0}$ if $U$ is a topological cap and $U\setminus V$ is an $\varepsilon$-neck, where $V$ is a compact domain in $U$ diffeomorphic to some $D^{n}//\Gamma$ or $\mathbb{S}^{n}//\langle\Gamma,\hat{\sigma}\rangle\setminus B$ or $\mathbb{S}^{n}//(x,\pm x^{\prime})\setminus B$ with $\partial V$ a central cross-section of an $\varepsilon$-neck (that is, the image of $\mathbb{S}^{n-1}/\Gamma\times\{0\}$ under the diffeomorphism associated to the $\varepsilon$-neck) and with $x_{0}\in\text{Int}\hskip 1.42262ptV$, (we will call $V$ a core of the cap $U$,) and in addition,
@@
The idea behind this definition is that an $\varepsilon$-cap $U$ should be, after suitable rescaling, $\eta$-close to the corresponding subset $W$ of an orbifold ancient $\kappa$-solution or an orbifold standard solution with an $\eta$-homothety $\psi:W\rightarrow U$, where $\eta$ is much smaller than $\varepsilon$ (compare p. 522 of [21] and Section 2.3 of [43]); moreover, the definition should help one to identify some components which are discarded in the process of the surgery of the Ricci flow (see Proposition 3.2). Assume that the end of the $\varepsilon$-cap $U$ is diffeomorphic to $\mathbb{S}^{n-1}\times(0,1)$. Let $N_{1}$ be an $\varepsilon_{0}$-neck given by $\psi_{1}:\mathbb{S}^{n-1}\times(-\varepsilon_{0}^{-1},\varepsilon_{0}^{-1})\rightarrow N_{1}$ which is contained in the $\varepsilon$-neck end in $U$. (Note that by Proposition A.4, given $0<\varepsilon<\varepsilon^{\prime}\leq 2\varepsilon_{1}$ and an $\varepsilon$-neck diffeomorphic to $\mathbb{S}^{n-1}\times(0,1)$, any $\varepsilon^{\prime}$-neck contained in this $\varepsilon$-neck is also diffeomorphic to $\mathbb{S}^{n-1}\times(0,1)$.) If $\varepsilon_{0}\leq\varepsilon_{4}$, by Lemmas A.2 and A.3, a suitable restriction of $\psi^{-1}\circ\psi_{1}$ will give a $2\varepsilon_{4}$-neck in $W$ if $\eta$ is sufficiently small. Then we can use Proposition 2.5 or Proposition 2.6 and the map $\psi$ to get the desired extra property for the $\varepsilon$-cap $U$.
@@
(cf. Theorem 6.18 in [21] and Theorem 3.9 in [31])   Given a small positive constant $\varepsilon$ and a constant $\theta>0$, there exist positive constants $C_{1}=C_{1}(n,\theta,\varepsilon)$ and $C_{2}=C_{2}(n,\theta,\varepsilon)$, such that given any noncompact orbifold ancient $\kappa$-solution $(\mathcal{O},g(t))$ of dimension $n\geq 5$ (with at most isolated singularities) which satisfies $Rm-\theta R\hskip 2.84526pt\text{id}\mathchoice{\mathbin{\ooalign{$\displaystyle\bigcirc$\cr$\displaystyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\textstyle\bigcirc$\cr$\textstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptstyle\bigcirc$\cr$\scriptstyle\wedge$\cr}}{}}{\mathbin{\ooalign{$\scriptscriptstyle\bigcirc$\cr$\scriptscriptstyle\wedge$\cr}}{}}\text{id}\in C_{\text{PIC}}$ and is not locally isometric to an evolving shrinking round cylinder, for each space-time point $(x_{0},t_{0})$, there is an open subset $U$ with $\overline{B(x_{0},t_{0},C_{1}^{-1}R(x_{0},t_{0})^{-\frac{1}{2}})}\subset U\subset B(x_{0},t_{0},C_{1}R(x_{0},t_{0})^{-\frac{1}{2}})$, which falls into one of the following categories:
(a) $U$ is an $\varepsilon$-neck diffeomorphic to $\mathbb{S}^{n-1}/\Gamma\times(-1,1)$ centered at $(x_{0},t_{0})$, or
(b) $U$ is an $\varepsilon$-cap diffeomorphic to $\mathbb{R}^{n}//\Gamma$ centered at $(x_{0},t_{0})$
@@
for some finite subgroup $\Gamma$ of $O(n)$ acting freely on $\mathbb{S}^{n-1}$; moreover, the scalar curvature in $U$ at time $t_{0}$ is (strictly) between $C_{2}^{-1}R(x_{0},t_{0})$ and $C_{2}R(x_{0},t_{0})$, and $vol_{g(t_{0})}(U)>(C_{2}|\Gamma|)^{-1}R(x_{0},t_{0})^{-n/2}$.